OpenStudy (anonymous):

Lowest common multiple of 3,2,& 5??

OpenStudy (jonnyvonny):

When in doubt, multiply them; 3*2*5= 6*5 = 30

OpenStudy (jonnyvonny):

To double check, select the most ugly one (3 in this case) and multiply till you get the answer you think it is, in this case, it is 30, so: 3,6,9,12,15,18,21,24,27,30. From these values, only 30 is divisible by both 2 and 5